THE ν_{5} BAND OF CH_{3}CD_{3}: HIGH RESOLUTION SPECTRUM AND GLOBAL THREE-BAND ANALYSIS

\underline{J.R. COOPER}, Department of Physics and Astronomy, University of Calgary, Calgary, Alberta T2N 1N4, Canada; A.R.W. McKELLAR, Steacie Institute for Molecular Sciences, National Research Council of Canada, Ottawa, Ontario K1A 0R6, Canada; I. OZIER, Department of Physics and Astronomy, University of British Columbia, Vancouver, B.C. V6T 1Z1, Canada; N. MOAZZEN-AHMADI, Department of Physics and Astronomy, University of Calgary, Calgary, Alberta T2N 1N4, Canada.

The lowest frequency parallel fundamental band ν_{5} of CH_{3}CD_{3} near 900 cm^{−1} was measured at low temperature with a resolution of 0.0021 cm^{−1} using Fourier transform spectroscopy. The band is weak, and an absorption path of 60m was used. Large torsional splittings due to inter-vibrational coupling have been observed. Building on previous studies of the torsional levels in the ground vibrational state and and in the methyl rocking state (ν_{12} = 1), a three-band analysis including this most recent data has been completed. The combined data set of more than 2,200 frequencies was fitted to within experimental accuracy using a 43-term model Hamiltonian. The results were found to bear a striking resemblance to those of an earlier, analogous study of CH_{3}SiH_{3}. In both cases, Fermi coupling between the (ν_{5} = 1) state and the ground state was found to be the dominant interaction responsible for the observed torsional splittings. Inclusion of this coupling results in a simplification of the ground-state Hamiltonian, so that only eight additional terms were required with the introduction of the ν_{5} band.